OpenStudy (anonymous):

Write an equation for teh line passing through (-11, -12) and parallel to x=2

OpenStudy (amistre64):

hmmmm

OpenStudy (katrinakaif):

An equation parallel would have the same slope. so in teh new equation: (y-y1) = m (x-x1) (y +12) = 2(x -11) and solve.

OpenStudy (angela210793):

parallel----> m=m1=2 y-y1=m(x-x1) y+12=2(x+11) y=2x+10

OpenStudy (amistre64):

we need a line that has no y values that hits x = -11 hmm

OpenStudy (amistre64):



OpenStudy (amistre64):

i cant see how you define the slope to be 2 ...